3x - 4y + 3 = 0
6x - 8y +7 = 0
a. Find the distance between the two lines.
b. Find the equation of the perpendicular line passing through. (-6,4)
c. Determine the distance of 2 given equations to point (-6.4)
1
Expert's answer
2020-05-08T19:17:01-0400

"d=\frac{|Ax_0+By_0+C|}{\sqrt{A^2+B^2}}" - formula to find distance between some point "M(x_0,y_0)" and a line "Ax+By+C=0"

We can use it to solve our first problem.


a)

Let "x=0", then "3\cdot 0-4y+3=0 \iff y=\frac{3}{4}"

So point "M(0,\frac{3}{4})" belong to the first line.

"A_2=6, B_2=-8, C_2=7" - coeffiсients of second line.

"d=\frac{|6\cdot 0+(-8)\cdot\frac{3}{4}+7|}{\sqrt{6^2+(-8)^2}}=\frac{|1|}{\sqrt{100}}=\frac{1}{10}"

(Lines don't intersect because "\frac{A_1}{A_2}=\frac{B_1}{B_2}\neq\frac{С_1}{С_2}". They are parallel)


b) If two lines are perpendicular, their normal vectors are perpendicular too.

"\vec {n_1}=\{3;-4\}" normal vector of first line.

"\vec {n_3}=\{a,b\}" normal vector of perpendicular line

"\vec{n_1}\cdot\vec{n_3}=3\cdot a+(-4)\cdot b=0" (scalar product )

Let "a=4," then "b=3"

So "\vec{n_3}=\{4,3\}" (what means "A_3=4, B_3=3" )

"4\cdot(-6)+3\cdot4+C_3=0" (substitution of (-6, 4))

"C_3=12"

"4x+3y+12=0" - perpendicular line


c) We can use our first formula

"d=\frac{|Ax_0+By_0+C|}{\sqrt{A^2+B^2}}"

First line:

"d_1=\frac{|3\cdot(-6)+(-4)\cdot4+3|}{\sqrt{3^2+(-4)^2}}=\frac{|-31|}{\sqrt{25}}=\frac{31}{5}=6.2"

"d_2=\frac{|6\cdot(-6)+(-8)\cdot4+7|}{\sqrt{6^2+(-8)^2}} =\frac{|-61|}{\sqrt{100}}=\frac{61}{10}=6.1"